Leather, Inc has two products and it's contribution margin income statement is as follows:

Product A Product B Total
Sales $100 $300 $400
Variable Costs $40 $200 $240
Contribution Margin $60 $100 $160
Fixed Costs $200
Net Loss $(40)

1.) Calculate the average contribution margin ratio for Leather, INC.

2.) Calculate the total sales dollar amount needed from both products for Leather, INC to break even.
Answer #1

1)

Contribution Margin of product A = $60

Contribution Margin of product B = $100

Sales of product A = $100

Sales of product B = $300

Total sales = $400

Contribution margin ratio, Product A = Contribution margin/Sales

= 60/100

= 60%

Contribution margin ratio, Product B = Contribution margin/Sales

= 100/300

= 33.3333333333%

Average contribution margin ratio = Contribution Margin of product A x Sales of product A/Total sales + Contribution Margin of product B x Sales of product B/Total sales

= 60% x 100/400 + 33.3333333333% x 300/400

= 15% + 25%

= 40%

2)

Break even sales in dollars = Fixed cost/Average Contribution margin ratio

= 200/40%

= $500

Please ask if you have any query related to the question. Thank you